# NoteLLM-2: Multimodal Large Representation Models for Recommendation

### v1

@@ 4 Experiments
  MTomato-Base 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
   + mICL 74.16 90.95 98.26 51.56 75.86 93.84 52.71 78.12 95.82
   + late fusion 74.51 91.37 98.45 54.19 78.02 94.50 54.00 79.71 95.80
   + NoteLLM-2 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
only late fusion 74.22 91.25 98.43 53.21 77.75 94.65 53.65 79.77 95.59
  MQwen-Base 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
   + mICL 75.82 91.57 98.55 53.32 76.40 94.38 54.52 79.31 96.05
   + late fusion 75.92 91.49 98.28 52.74 76.03 93.65 54.06 78.89 95.53
   + NoteLLM-2 76.50 92.18 98.58 55.72 78.91 94.94 56.13 80.65 96.24
only late fusion 75.84 91.49 98.35 53.36 76.23 93.57 54.27 78.39 95.30
  MQwen-bigG 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
   + mICL 77.97 93.19 98.92 58.50 82.08 96.66 58.58 84.01 97.20
   + late fusion 77.43 92.92 98.78 57.47 81.77 95.89 58.50 83.99 97.16
   + NoteLLM-2 77.56 93.38 98.82 57.94 81.98 96.29 58.79 83.87 97.05
only late fusion 76.24 92.31 98.67 55.21 79.22 95.58 55.15 81.13 96.70
4.1 Performance Evaluation
We conduct experiments on all components of NoteLLM-2 based on three MLRMs to show our method effectiveness. Besides, we perform an ablation study, only late fusion, which only use late fusion to integrate image and text information, without inputting image embeddings into LLMs.
The results are presented in Table 3. Firstly, NoteLLM-2 can significantly enhance the overall performance for MTomato-Base and MQwen-Base, which have relatively small vision encoders. Additionally, NoteLLM-2 primarily enhances MQwen-bigG in terms of short pairs. Secondly, mICL enhances the performance of all models, while late fusion is more effective in models with relatively small vision encoders. Concurrently, we find that only late fusion is a straightforward and efficient fusion method. However, its simplicity could potentially be a drawback when the vision encoder is more powerful, as it may not fully and effectively interact with LLMs, leading to a loss in performance. Lastly, MQwen-bigG + NoteLLM-2 still underperforms compared to Qwen-VL-Chat, especially in shorter pairs. We attribute this to the significant difference in visual token length between MQwen-bigG ($16$) and Qwen-VL-Chat ($256$).
((a)) MTomato-Base
((b)) MQwen-Base
((c)) MQwen-bigG
Figure 4: Relative sizes of saliency scores in different layers of different MLRMs.
4.2 Saliency Scores of Enhanced MLRMs
To further understand the impact of NoteLLM-2 on MLRMs, we illustrate the difference in saliency scores of enhanced MLRMs. The saliency scores for the original fine-tuned method are $S_{v}$, $S_{t}$ and $S_{o}$. The saliency scores for the enhanced fine-tuned method are $\hat{S_{v}}$, $\hat{S_{t}}$ and $\hat{S_{o}}$. We consider the visual note compressed word as a part of the vision embeddings $E_{v}$. Our results are shown in Figure 4. The representations of all enhanced MLRMs amplify the direct focus on images while reducing the emphasis on $S_{o}$ in the shallow layers. At the same time, $S_{t}$ remains virtually unchanged. This is attributed to mICL’s use of the same compressed prompt to compress both modalities. By identifying similar compression patterns from image information, mICL can bolster the concentration of multimodal representations on images, similar to ICL [48].
Table 4: Impact of the number of visual tokens $L$ (%).
  #Visual Tokens All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  8 74.17 90.85 98.16 52.80 76.13 93.76 53.19 77.87 94.48
16 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
32 74.55 91.41 98.38 54.56 78.91 94.94 54.21 79.83 95.63
48 74.30 91.57 98.48 54.31 78.83 95.13 53.73 80.06 95.92
Table 5: Impact of textual loss and visual loss ratio $\alpha$ (%).
@@ 4.2 Saliency Scores of Enhanced MLRMs
4.3 Hyper-Parameter Analysis
In this section, we use MTomato-Base to conduct hyper-parameter analysis experiments.
Impact of length of visual tokens. We explore the impact of visual token length. The results are shown in Table 4. Reducing the length from $16$ to $8$ worsens performance in short pairs. There exists a tradeoff between performance and inference speed concerning the length of visual tokens.
5 Conclusion
@@
This paper explores using LLMs to improve textual comprehension in multimodal representation tasks under I2I recommendation scenarios. We design an end-to-end training method that can customize the integration of any existing LLMs and vision encoders. This approach reduces reliance on open-source MLLMs, which require costly multimodal pre-training. To prevent the problem of ignoring visual information in end-to-end training, we propose NoteLLM-2, which contains mICL and late fusion. The effectiveness of this approach is confirmed through extensive experiments.
@@
Appendix A Related Work
A.1 Multimodal Representation
With the advancement of pre-training techniques, numerous studies utilize web-scale image-text data to pre-train models for a fundamental understanding of multimodal information. There are two main pre-training techniques. The first technique adopts generative pre-training objectives, following masked language/image modeling [14, 49, 27] or image captioning [56, 26]. The second technique uses contrastive pre-training objectives, such as contrastive learning [37, 58, 45, 46, 56, 49] and image-text matching tasks [14, 27, 26]. These models demonstrate a powerful zero-shot ability in many downstream tasks. However, the text comprehension capabilities of these models are unsatisfactory [6, 59], particularly in the era of LLMs. In this work, we explore to use LLMs to enhance the multimodal representations.
A.2 Large Language Models
LLMs have attracted extensive attention due to their remarkable text understanding and generation capability [18, 7, 2, 47]. Therefore, numerous text applications have been developed based on LLMs [31, 60, 34, 52, 22]. To understand more modality information, LLMs are extended into MLLMs [3, 5, 55, 33, 24]. Most MLLMs employ specialized modality encoders to capture non-textual information [42, 3, 5]. This information is then mapped into textual LLM space using connectors and subsequently processed by LLMs. Due to the significant understanding ability of LLMs, MLLMs can generate answers according to textual instruction and multimodal information [15, 13, 3, 5]. While existing studies primarily focus on the generation abilities of LLMs, some explore the text representation ability of LLMs [34, 41, 60, 19, 31, 4], demonstrating significant potential to enhance performance for numerous downstream tasks. However, no studies have explored the multimodal representation ability with the assistance of LLMs. Therefore, we aim to investigate the multimodal representation ability enhanced by LLMs in recommendation scenarios.
A.3 Multimodal I2I Recommendation
I2I recommendation generates a ranked list of items from a vast pool based on a specific target item, which is an important technique for recall stages in recommendation [54, 64, 60]. I2I recommendation employs either a pre-constructed I2I index [54] or the approximate k-nearest neighbor method [21] to retrieve relevant items online. Some studies have investigated multimodal content-based I2I recommendations [57, 10, 20, 63, 16]. For example, CommerceMM [57] adopts multimodal pre-training tasks used in METER [14] and omni-retrieval tasks to train commerce multimodal representation. MAPS [10] proposes a multimodal attention framework to measure product similarity. However, these works adopt small-scale pre-trained visual-linguistic models as backbones, which have huge potential for scaling up. As shown in [60, 19, 4], LLMs can outperform smaller models in textual representation tasks significantly. Therefore, in this work, we explore empowering the multimodal representations with LLMs in multimodal I2I recommendation.
Appendix B Limitations and Future Works
More modalities. In this work, our focus is primarily on the processing of multimodal situations that involve the input of a single image and corresponding text information. This is the most simplified situation within our research scenarios. In fact, the scope of our representation method can extend to include audio and video data [42, 32, 28]. This direction presents an opportunity for further exploration and expansion into more complex modalities.
General scenarios. Our MLRMs mainly focus on the multimodal representation ability in recommendation scenarios. These scenarios determine item similarity by analyzing the interaction between items and users. However, this is not a universally applicable scenario. The current focus of our MLRMs is somewhat specialized and may not cover all potential use cases. Given the potential of MLRMs, it is promising to consider broadening their multimodal representation ability into more general scenarios [35].
Larger and stronger base models. Due to the rapid development of LLMs and vision encoders, there are an extensive range of larger and more powerful base models available in the market [58, 2, 11]. However, due to resource limitations and the demand for rapid recommendations, we have not yet experimented with all of these models. There is still a vast scope to investigate more advanced models, which could potentially boost performance even further.
Appendix C Dataset Details
Table 6: Detailed statistics of training and testing dataset.
                                  training dataset
  # notes 1,487,613 # note pairs 1,130,906
avg. # words per title 12.38 avg. # topic per note 5.26
avg. # words per topic 4.53 avg. # words per content 133.20
                                  testing dataset
  # notes 534,144 # note pairs 21,262
avg. # words per title 11.68 avg. # topic per note 4.58
avg. # words per topic 4.50 avg. # words per content 116.00
Appendix D Training Details
Table 7: Detailed training hyperparameters for training MLRMs.
@@ Appendix D Training Details
For all models with LLMs, we list the detailed training hyperparameter in Table 7. Besides, we use DeepSpeed [38], and Zero Redundancy Optimizer (ZeRO) [39] Stage 3 to train our models.
Appendix E Model Details
E.1 Baseline Details
To evaluate the effectiveness of MLRMs, we compare them with several baselines. (1) CLIP ViT-B [37] is trained using only image input. (2) BM25 [40] is a basic baseline, which counts the relevance between the query and documents based on every word in the query. (3) RoBERTa-wwm-ext [9] is a classic traditional Chinese BERT-like [23] text encoder with pure text input. (4) Tomato is our continually pre-trained LLM that is based on LLaMA 2 [47] and utilizes data from our platform for training. (5) NoteLLM [60] unifies the embedding generation and topic generation into one task. (6) METER Merge-attn and METER Co-attn [14] leverage best training strategy to train vision-and-language transformers. METER Merge-attn adopts the merged attention as the fusion module, while METER Co-attn uses co-attention as the fusion module. All baselines are trained on our training dataset.
E.2 End-to-End MLRM Details
We provide the details in Table 8.
Table 8: Model settings of end-to-end MLRMs.
@@ E.2 End-to-End MLRM Details
E.3 Comparison of MLRMs
We provide the comparison in Table 9.
Table 9: Comparison of different LLMs for representation tasks. $L_{c}$ means the number of vision embeddings for each image. The inference speed test is conducted on an A100 GPU, with an inference batch size set at $50$. All notes in this test are the longest note in the dataset. The speed is quantified as the number of longest notes inferred per second.
  Method Vision Encoder LLM VL Connector # Total Par. $L_{c}$ Inf. Speed
  Tomato - Tomato (6.7B) - 6.7B - 59.0
Qwen-Chat - Qwen-Chat (7.1B) - 7.1B - 57.0
  BLIP-2 ViT-g (1.0B) OPT (6.7B) Q-Former (105.1M) 7.8B 32 44.2
LLaVA-1.5 ViT-L (303.5M) Vicuna (6.7B) MLP (21.0M) 7.1B 576 12.1
Qwen-VL-Chat ViT-bigG (1.8B) Qwen-Chat (7.1B) Resamplers (76.1M) 9.0B 256 11.5
MTomato-Base ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.8B 16 51.5
    + NoteLLM-2 ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.9B 16 48.2
MQwen-Base ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 50.5
    + NoteLLM-2 ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 46.9
MQwen-bigG ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.1B 16 36.8
    + NoteLLM-2 ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.2B 16 35.2
Appendix F More Information about Saliency Scores
F.1 Formal Definition of $S_{v}$, $S_{t}$ and $S_{o}$
$S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$:
$S_{v}$ $=\frac{\sum_{(i,j)\in P_{v}}I_{l}(i,j)}{\left|P_{v}\right|},$ (9)
$P_{v}$ =$=\{(c,j):j\in E_{v}\}.$
$S_{t}$ $=\frac{\sum_{(i,j)\in P_{t}}I_{l}(i,j)}{\left|P_{t}\right|},$ (10)
$P_{t}$ =$=\{(c,j):j\in E_{t},j<c\}.$
$S_{o}$ $=\frac{\sum_{(i,j)\in P_{o}}I_{l}(i,j)}{\left|P_{o}\right|},$ (11)
$P_{o}$ =$=\{(i,j):j<i\}-P_{v}-P_{t}.$
F.2 Results of Qwen-VL-Chat
Figure 5: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of Qwen-VL-Chat.
We observe a unique pattern of saliency scores in Qwen-VL-Chat, as shown in Figure 5. The predominant information flow in Qwen-VL-Chat is textual, with a value higher than $0.8$ in all layers. This result is attributed to the extensive visual embeddings ($256$ in Qwen-VL-Chat) inputted into LLMs, which reduces the average saliency of image information flow. This denotes that traditional MLLMs are not directly suitable for MLRMs. The necessary action to adapt MLLMs to MLRMs is enhancing the efficiency of information flow.
Appendix G Performance of Zero-shot Multimodal Representation
Table 10: Zero-shot representation performance of different models (%).
@@ Appendix G Performance of Zero-shot Multimodal Representation
BM25 Text 55.20 72.67 82.83 35.82 48.42 59.71 35.07 49.63 63.68
In this section, we explore the multimodal representation ability of MLLMs in a zero-shot way in the multimodal I2I recommendation task. We select four popular MLLMs, i.e., BLIP-2 [25], LLaVA-1.5 [29], Qwen-VL [3] and Qwen-VL-Chat [3]. We also compare some baselines BM25 [40], Qwen-Chat [2] and our continual pre-trained LLM Tomato, which is pre-trained on our platform data based on LLaMA 2 [47], but lacks vision perception ability. The MLLMs details are concluded in Table 9. To understand the representation ability of MLLMs for different modalities, we also input images and texts independently for the representation test.

### v2

@@ 4.3. Hyper-Parameter Analysis
  8 51.0 74.17 90.85 98.16 52.80 76.13 93.76 53.19 77.87 94.48
16 48.2 74.62 91.50 98.47 54.87 78.82 94.69 54.51 79.75 95.46
32 44.2 74.55 91.41 98.38 54.56 78.91 94.94 54.21 79.83 95.63
48 41.0 74.30 91.57 98.48 54.31 78.83 95.13 53.73 80.06 95.92
Table 6. Impact of textual loss and visual loss ratio $\alpha$ (%).
@@
Figure 9. The online serving pipeline includes offline and online phases. The offline phase generates embeddings with the MLRM trained with NoteLLM-2 and updates the ANN index with new notes. The online phase recalls notes from the index based on the user’s interaction history.
Table 7. Comparison of different LLMs and MLRMs for representation tasks. $L_{c}$ means the number of vision embeddings for each image. All models are trained on A100 GPUs. The training cost is measured in GPU hours (GPU-h). The inference speed test is conducted on an A100 GPU, with an inference batch size set at $50$. All notes in this test are the longest note in the dataset. The speed is quantified as the number of longest notes inferred per second.
  Method Input Vision Encoder LLM VL Connector # Total Par. $L_{c}$ Training GPU-h Inf. Speed
  Tomato Text - Tomato (6.7B) - 6.7B - 78.3 59.0
Qwen-Chat Text - Qwen-Chat (7.1B) - 7.1B - 86.0 57.0
  BLIP-2 Multimodal ViT-g (1.0B) OPT (6.7B) Q-Former (105.1M) 7.8B 32 110.8 44.2
Qwen-VL-Chat Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Resamplers (76.1M) 9.0B 256 432.6 11.5
  MTomato-Base Multimodal ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.8B 16 108.6 51.5
    + NoteLLM-2 Multimodal ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.9B 16 110.7 48.2
  MQwen-Base Multimodal ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 100.3 50.5
    + NoteLLM-2 Multimodal ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 101.6 46.9
  MQwen-bigG Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.1B 16 129.8 36.8
    + NoteLLM-2 Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.2B 16 123.1 35.2
4.4. Efficiency Analysis
We provide a comparison of the model parameter size, training cost, and inference speed of LLMs in Table 7. State-of-the-art MLLMs such as Qwen-VL-Chat, designed for multimodal understanding and answering, have high training costs for adaptation and slow inference speed, making them impractical for fast-paced recommendation tasks. On the other hand, MLRMs designed for representation tasks are more efficient for both training and inference. Additionally, due to NoteLLM-2’s longer prompt template, integrating NoteLLM-2 usually slightly slows down training and inference, but achieves better performance.
4.5. Online Serving and Experiments
We present our online serving pipeline in Figure 9, which consists of offline and online phases. During the offline phase, the MLRM trained with NoteLLM-2 generates multimodal embeddings in real-time for newly generated notes from users. These embeddings are saved in an embedding table for reuse. Furthermore, the I2I ANN index is dynamically updated with embeddings of new notes. In the online phase, the system selects history embeddings from the existing embedding table based on the user’s interaction history notes. Using these history note embeddings, the system performs multiple ANN searches to find relevant notes in the I2I index. The top-k relevant notes are recalled and passed to the next recommendation stage. Now, our model has replaced the previous I2I recall process and has already become a crucial recall channel on the platform.
We conduct week-long online experiments on our platform by allocating 10% of the overall traffic for A/B testing. The online baseline uses the same LLM as our online method for I2I training, with the training approach utilizing NoteLLM (Zhang et al., 2024a). Our MLRM trained with NoteLLM-2 enhances the click count of the first thousand exposures by $6.35\%$ and the number of interactions within the first $24$ hours of note publication by $8.08\%$. Besides, notes with the first interaction within 24 hours and 100 exposures increase by $9.83\%$.
5. Related Work
Multimodal Representation. With the advancement of pre-training techniques, numerous studies utilize web-scale image-text data to pre-train models for a fundamental understanding of multimodal information. There are two main pre-training techniques. The first technique adopts generative pre-training objectives, following masked language/image modeling (Dou et al., 2022; Wang et al., 2023a; Li et al., 2021) or image captioning (Yu et al., 2022b; Li et al., 2022). The second technique uses contrastive pre-training objectives, such as contrastive learning (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023, 2024; Yu et al., 2022b; Wang et al., 2023a) and image-text matching tasks (Dou et al., 2022; Li et al., 2021, 2022). These models demonstrate a powerful zero-shot ability in many downstream tasks. However, the text comprehension capabilities of these models are unsatisfactory (Chen et al., 2023; Zhang et al., 2024c), particularly in the era of LLMs. In this work, we explore the use of LLMs’ powerful comprehension capabilities to enhance multimodal representations.
Large Language Models. LLMs have attracted extensive attention due to their remarkable text understanding and generation capability (Jiang et al., 2023b; Chiang et al., 2023; Bai et al., 2023a; Touvron et al., 2023). Therefore, numerous text applications have been developed based on LLMs (Ma et al., 2023; Zhang et al., 2024a; Muennighoff et al., 2024; Xu et al., 2024; Karamcheti et al., 2024). To understand more modality information, LLMs are extended into MLLMs (Bai et al., 2023b; Chen et al., 2024; Yin et al., 2023; McKinzie et al., 2024; Laurençon et al., 2024). Most MLLMs employ specialized modality encoders to capture non-textual information (Su et al., 2023; Bai et al., 2023b; Chen et al., 2024). This information is mapped into textual space and processed by LLMs, enabling MLLMs to generate answers from multimodal input (Ge et al., 2023; Dong et al., 2023; Bai et al., 2023b; Chen et al., 2024). While existing studies primarily focus on the generation abilities of LLMs, some explore the text representation ability of LLMs (Muennighoff et al., 2024; Springer et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023; BehnamGhader et al., 2024). However, no studies have explored multimodal representation with LLMs. Therefore, we investigate the multimodal representation ability enhanced by LLMs in recommendation scenarios.
Multimodal I2I Recommendation. I2I recommendation generates a ranked list of items from a vast pool based on item similarity, which is an important technique for recall stages in recommendation (Yang et al., 2020; Zhu et al., 2018; Zhang et al., 2024a). I2I recommendation employs a pre-constructed I2I index (Yang et al., 2020) to retrieve relevant items by the approximate k-nearest neighbor method (Johnson et al., 2019). Some studies have investigated multimodal content-based I2I recommendations (Yu et al., 2022a; Das et al., 2022; Jin et al., 2023; Zheng et al., 2023; Han et al., 2023). However, these works adopt small-scale pre-trained visual-linguistic models as backbones, which have huge potential for scaling up. Therefore, in this work, we explore empowering the multimodal representations with LLMs in multimodal I2I recommendations.
6. Conclusion and Future Work
@@
This paper explores using LLMs to improve textual comprehension in multimodal representation tasks under I2I recommendation scenarios. We design an end-to-end fine-tuning method that can customize the integration of any existing LLMs and vision encoders. This approach reduces reliance on open-source MLLMs, which require costly multimodal pre-training. To prevent the problem of ignoring visual information in end-to-end fine-tuning, we propose NoteLLM-2, which contains mICL and late fusion. The effectiveness of this approach is confirmed through extensive experiments. In future work, we will extend our method to more complex modalities, like videos, and broader scenarios and recommendation tasks.
Acknowledgements
This work was supported in part by the grants from National Science and Technology Major Project (No. 2023ZD0121104), and National Natural Science Foundation of China (No.62222213, 62072423). Xiangyu Zhao was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), CCF-BaiChuan-Ebtech Foundation Model Fund, Collaborative Research Fund No.C1043-24GF, and Kuaishou.
@@ Acknowledgements
Appendix A Training Details
Table 8. Detailed training hyperparameters for fine-tuning MLRMs.
@@ Appendix A Training Details
For all models with LLMs, we list the detailed training hyperparameter in Table 8. Besides, we use DeepSpeed (Rasley et al., 2020), and Zero Redundancy Optimizer (ZeRO) (Ren et al., 2021) Stage 3 to train our models.
Appendix B End-to-End MLRM Details
We provide the details about end-to-end MLRMs in Table 9. Besides, we present the continuous pre-training details of Tomato to enhance understanding of its characteristics. Tomato is based on LLaMA 2 (Touvron et al., 2023), which originally only supports English. However, since our platform targets the Chinese market, it is necessary to enhance LLaMA 2’s Chinese understanding capabilities. Thus, we have made the following improvements to continuous pre-training:
  • Vocabulary Expansion: We extended the original vocabulary from $32,000$ tokens to $49,216$ tokens. The additional $17,216$ tokens are primarily Chinese.
  • Pre-training Data: We collect pre-training data from three sources: Chinese corpora, English corpora, and Platform-specific Data. We utilize open-source Chinese corpora, including Wanjuan (He et al., 2023) and WuDao (Xue et al., 2022), totaling approximately 2T tokens. To maintain English proficiency, we also include RedPajama (Computer, 2023). Besides, we incorporate 30B tokens of high-quality notes collected from our platform to better understand the data characteristics of our platform.
Tomato enhances Chinese understanding while maintaining its English capabilities at a level comparable to the original LLaMA 2 model. It is worth noting that the pre-training details are unrelated to our contributions. Our method is based on the model after continual pre-training. This also shows an advantage of our method, which allows for the personalized enhancement of the base models for different modalities before fine-tuning using NoteLLM-2.
Table 9. Model settings of end-to-end MLRMs.
@@ Appendix B End-to-End MLRM Details
Table 10. Zero-shot representation performance of different models under different modal inputs (%). “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
@@
BM25 Text 55.22* 72.64* 82.82* 35.75* 48.41* 59.81 35.05* 49.52* 63.64
Appendix C Saliency Scores of Qwen-VL-Chat
We observe a unique pattern of saliency scores in Qwen-VL-Chat, as shown in Figure 10. The predominant information flow in Qwen-VL-Chat is textual, with a value higher than $0.8$ in all layers. This result is attributed to the extensive visual embeddings ($256$ in Qwen-VL-Chat) inputted into LLMs, which reduces the average saliency of image information flow. This denotes that traditional MLLMs are not directly suitable for MLRMs. The necessary action to adapt MLLMs to MLRMs is to increase the density of information flow.
Figure 10. Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of Qwen-VL-Chat.
Appendix D Performance of Zero-shot Multimodal Representation
In this section, we explore the multimodal representation ability of MLLMs in a zero-shot way in the multimodal I2I recommendation task. We select four popular MLLMs, i.e., BLIP-2 (Li et al., 2023), LLaVA-1.5 (Liu et al., 2023b), Qwen-VL (Bai et al., 2023b) and Qwen-VL-Chat (Bai et al., 2023b). We also compare some baselines BM25 (Robertson et al., 2009), Qwen-Chat (Bai et al., 2023a) and our continual pre-trained LLM Tomato, which is pre-trained on our platform data based on LLaMA 2 (Touvron et al., 2023), but lacks vision perception ability. The MLLMs details are concluded in Table 7. To understand the representation ability of MLLMs for different modalities, we also input images and texts independently for the representation test.